OpenStudy (anonymous):

Help with this Find the equation of the tangent line y=tanx at (pi/4,1) verify graphically

OpenStudy (campbell_st):

find the 1st derivative, that gives you the equation of the slope then substitute pi/4 to find the slope. Now you can find the equation, you have the slope and point hope it makes sense.

OpenStudy (anonymous):

so the first derivative is -sinx

OpenStudy (anonymous):

wait no. My bad i meant it's sec^2x right?

OpenStudy (campbell_st):

yes so substitute x = pi/4 to find the slope sec = 1/cos so the slope is \[\sec^2(\frac{\pi}{4})=(\sec(\frac{\pi}{4}) )^2= (1/\cos(\pi/4))^2 = (\frac{\sqrt{2}}{1})^2\]

OpenStudy (anonymous):

okay i understand everything except how u got (sqrt2/1)^2

OpenStudy (campbell_st):

ok you are working with an exact value  \[\cos(\frac{\pi}{4}) = \frac{1}{\sqrt{2}}\] since sec = 1/cos take the reciprocal of the exact value of sec(pi/4) \[\sec(\frac{\pi}{4})= \frac{1}{\cos(\frac{\pi}{4})} = \frac{1}{\frac{1}{\sqrt{2}}} = \sqrt{2} \]